OpenStudy (anonymous):

Okay, I'm really confused on this one: What is the equation of a line, in general form, that passes through points (-1, 2) and (5, 2)? y - 2 = 0 x - 2 = 0 y - x - 2 = 0

OpenStudy (anonymous):

@sleepyjess Stuart, can you help please c:

OpenStudy (anonymous):

m = (y2-y1)/(x2-x1) general form y-y1 = m(x-x1) Put the given values in respective terms

OpenStudy (anonymous):

General form is Ax+By-C=0. The slope I got was 6/0

OpenStudy (anonymous):

m = 2-2/5-(-1) m = 0/6 m = 0 y-2 = 0*(x-5) y -2 = 0

OpenStudy (anonymous):

what

OpenStudy (anonymous):

@Sparklegirl14 whats the confusion?